OpenStudy (anonymous):

The drama dance sold 1500 tickets for the end of the year performance. Admission prices were $12 for adults and $6 for students. The total amount collected at the box office was $16,200. How many students attended the play? Help.

OpenStudy (anonymous):

Try makeing a set of functions: 1500 = a + s 16,200 = 12a + 6s So now use the substitution method: 1500 - a = s 16,200 = 12a + 6(1500 - a) = 12a + 9000 - 6a 16,200 = 6a + 9000 -9000 7200/6 = 6a/6 1200 = a Substitute into original problem: 1500 = 1200 + s So how many students went?

OpenStudy (anonymous):

300 thank you.

OpenStudy (anonymous):

No problem! :)